The table below lists the prices and quentrics consumed af thrce different goads fram 2010-2012 2010 2011 2012 ntity n For 2010 2011, and 2012 determine the nmaunt thnt 츠 Typical consumer pays encn yeerto purchase the quantities listed in the table above Instructions: Round your answers to the nearest whole number. 2011 2012 aurner expencitureS Instructions: Round your answers to wo decimal places. b The pe centage cnonge in the ormount the consumer paid is[ %from 200 to 2011, and -from 2011 to 2012. c. It Is problematic to use your answers to port b as a measure of Inflation becouse Clck to select) Instructions: Round your answers to two decimal places. d. Suppose we teke 2010 as the base year, implying that the market besket is fixed at 2010 consumption levels. Usng 2010 consumption levels, the rate of inlation ton, 2010 to 2on, andL %from 2011 to 2012. Round your ansers ta two decimal places e. Repeat th exercise f om part d nowassuming that the base year is 2011 Usng 2011 con un pion levels the tate of rf a on s % om 2010 to 2011 nd % from 2011 to 2012. . Your aswers from parts d and e were dilferet because(Click lu select ick to select pricss have changad ircome has changed he base years put different weight on the goud he base years have the same consumption quentties References eBook &Resources
Answer #1

a.

2010

2011

2012

Consumer Expenditure

196.0

180.0

260.0

Consumer Expenditure as 2010 base year

196.0

215.0

147.0

Consumer Expenditure as 2011 base year

202.0

180.0

175.0

b. 2010 to 2011 =((180-196)/196)*100 =-8.2%
2011 to 2012 =((260-180)/180)*100 =44.4%

c. It shows the changes in expenditure level

d. 2010 to 2011 =((215-196)/196)*100 9.7%
2011 to 2012 =((147-215)/215)*100 =-31.6%

e. 2010 to 2011 =((180-202)/202)*100 = -10.9%
2011 to 2012 =((175-180)/180)*100 = -2.8%

f. The base year puts different weights on goods